OpenStudy (anonymous):

determine sin(2x) when cos(x)=-√(3)/2 and sin(x) is positive?

OpenStudy (ajprincess):

\(\sin(2x)=2\sin x\cos x\) \(\sin x=\sqrt{1-\cos^2x}\) Does that help? @jsaavedr